mese48:

Which does the ratio 17:44 represent?

mese48:

After-School Activities Activity Number of Students Spanish Club 19 Basketball 24 Drama Club 15 Homework Help 17 Soccer 20 Which does the ratio 17:44 represent? the part-to-part relationship between homework help and sports the part-to-whole relationship between homework help and sports the part-to-part relationship between homework help and all activities the part-to-whole relationship between homework help and all activities

jordangantt12:

Sorry for the late reply, but as you see 17 represents homework help. If you add up the numbers from all the groups, you get 44. Therefore it would be a part to whole relationship between homework help and all activities.

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jordangantt12 Sorry for the late reply, but as you see 17 represents homework help. If you add up the numbers from all the groups, you get 44. Therefore it would be a part to whole relationship between homework help and all activities. \(\color{#0cbb34}{\text{End of Quote}}\) please explain it how you get this 44 ? thank you in advance

jhonyy9:

without this 17 if you add up the numbers how you get 44 ?

jhonyy9:

44 you can get for example if you add up 20 and 24

jhonyy9:

and there are again 15 and 19

jordangantt12:

Yes I'm sorry . I was in between two questions and ended up putting both answers together .

jhonyy9:

but 15 +19 = 34 not 44

jhonyy9:

@jordangantt12 i accept your answer but please explain me how you get this relation and what choice id correct ?

jordangantt12:

yes 20 +24 = 44 . 20 & 24 are sports . Therefore it would have to be the part-to-whole relationship between homework help and sports

jordangantt12:

so 17:44 ration is homework help to sports .

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jordangantt12 Sorry for the late reply, but as you see 17 represents homework help. If you add up the numbers from all the groups, you get 44. Therefore it would be a part to whole relationship between homework help and all activities. \(\color{#0cbb34}{\text{End of Quote}}\) ok but here you ve wrote in other way

jordangantt12:

yes I realized that . thank you for correctin me .

jhonyy9:

np yw but please more ATTENTION IN FUTURE !!!

jhonyy9:

now i see that who asked this problem is offline so he or she go away with bad answer

jhonyy9:

@jordangantt12 please more attention - ok ?

jhonyy9:

thank you and yw

jhonyy9:

@Laylalyssa any idea about this pls ? ty

jordangantt12:

yw and sorry .

jhonyy9:

ok hnd

Laylalyssa:

So it would be the part-to-whole relationship between homework help and sports because homework help : 17 Sports has basketball 24 and soccer 20 24+20=44 17:44 right?

jordangantt12:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa So it would be the part-to-whole relationship between homework help and sports because homework help : 17 Sports has basketball 24 and soccer 20 24+20=44 17:44 right? \(\color{#0cbb34}{\text{End of Quote}}\) ye .

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa So it would be the part-to-whole relationship between homework help and sports because homework help : 17 Sports has basketball 24 and soccer 20 24+20=44 17:44 right? \(\color{#0cbb34}{\text{End of Quote}}\) yes